OpenStudy (anonymous):

Write the slope-intercept equation for the line containing the given pair of points.(-2,0) and (0,3)

OpenStudy (hoblos):

slope = 3/2 y= 3/2 x +3

OpenStudy (anonymous):

I am sorry for asking these questions but I just don't get this.

OpenStudy (hoblos):

do you know how to get the slope?

OpenStudy (anonymous):

no not really.

OpenStudy (sasogeek):

difference in y components / difference in corresponding x components

OpenStudy (sasogeek):

rise/run

OpenStudy (sasogeek):

y2-y1/x2-x1

OpenStudy (hoblos):

yup exactly as sasogeek said

OpenStudy (anonymous):

Thank you both, I am going back to my textbook to get more of a understanding. Thx for your help.

OpenStudy (sasogeek):

i hope i made sense to you though, that's more important than the thank you :) but i appreciate it, let us know if you have any more questions, we're here for you :)

OpenStudy (anonymous):

Ok I will just do that.And yes it does make alot of sense to me.